Shown below are three scenarios in which a truck i
Answer #1

Suppose the force that car exert on boat trailer be the reation= Fc-b = F0 i^

Then the boat trailer according to newtons 3rd law will also exert the same force but in opposite direction

Thus, the force that boat trailer exert on car = Fb-c = Fc-b =F0

DIfference of strength of these two forces are always zero as shown, Fb-c -Fc-b = F0-F0 =0

Difference is zero for all the cases.